# Unitary discriminants of characters

### v1

@@ 7.1 The fixed algebra under an automorphism
Let $G$ be a finite group and $\chi\in\operatorname{Irr}^{o}(G)$. Put $2m:=\chi(1)$, $L={\mathbb{Q}}(\chi)$, and assume that there is an $L$-representation $\rho:G\to\operatorname{GL}(V)$ affording the character $\chi$. Put $K:=L^{+}$ to denote the real subfield of $L$ and let $H:V\times V\to L$ be a $\rho(G)$-invariant $L/K$-Hermitian form on $V$. Assume that there is $\alpha\in\operatorname{Aut}(G)$ with $\alpha^{2}=1$ such that $\chi\circ\alpha=\overline{\chi}$. Then $\chi+\overline{\chi}$ extends to an absolutely irreducible character
$\operatorname{{\mathfrak{X}}}={\textrm{Ind}}_{G}^{\mathcal{G}}(\chi)$
of the semidirect product ${\mathcal{G}}:=G:\langle\alpha\rangle$.
@@
    Then $[A]=[{\mathcal{Q}}]\in\operatorname{Br}_{2}(L,K)$.
@@
    Assume that the Frobenius-Schur indicator of $\operatorname{{\mathfrak{X}}}$ is $+$.
    Then $A$ is an orthogonal subalgebra of $(\operatorname{End}_{L}(V),\iota_{H})$.
@@
    Assume that the Frobenius-Schur indicator of $\operatorname{{\mathfrak{X}}}$ is --.
    Then $\operatorname{disc}(H)=\operatorname{disc}_{L}({\mathcal{Q}})^{m}$ and $\Delta(H)=[{\mathcal{Q}}]^{m}$.
@@
To identify the class of the central simple $K$-algebra $A$ in the Brauer group, we relate the centraliser of $A$ in $\operatorname{End}_{K}(V)$ to the Schur indices of $\operatorname{{\mathfrak{X}}}$, i.e. to the class $[{\mathcal{Q}}]\in\operatorname{Br}_{2}(L,K)$. We closely follow [2, Section 2]:
@@
Then $\overline{X}X$ commutes with all $\rho(g)$, so $\overline{X}X=\lambda I_{2m}$ for some $\lambda\in L^{\times}$. It is easy to see that $\lambda=\overline{\lambda}\in K^{\times}$ and is well defined up to norms. Now the proof of [2, Proposition 4.2] shows that $[{\mathcal{Q}}]=[(L,\lambda)_{K}]$.
@@
(c) It remains to show that the restriction of $\iota_{H}$ to $A$ is an orthogonal involution on $A$, i.e. that the dimension of skew-adjoint elements in $A$ is $m(2m-1)$. This follows from the assumption on the indicator of $\operatorname{{\mathfrak{X}}}$. As dimensions remain unchanged under field extensions we may replace $K$ by the field ${\mathbb{R}}$ of real numbers and $L$ by ${\mathbb{C}}$. The assumption on the Frobenius-Schur indicator of $\operatorname{{\mathfrak{X}}}$ shows that the Schur index of $\operatorname{{\mathfrak{X}}}$ over the real numbers ${\mathbb{R}}$ is 1. So the main result of [2] shows that there is a basis of $V\otimes_{L}{\mathbb{C}}$ such that the matrices in $\rho(G)$ with respect to this basis satisfy
@@
(d) Now assume that the indicator of $\operatorname{{\mathfrak{X}}}$ is --. Then by (a) the algebra $A\cong{\mathcal{Q}}^{m\times m}$ and the restriction of $\iota$ to $A$ is a symplectic involution. As in Corollary 4.10 one gets that $\Delta(H)=[{\mathcal{Q}}]^{m}$. ∎
@@
Let Weierstrass-$\wp$ be a prime ideal of $K$ that does not divide $2|G|$. Then Weierstrass-$\wp$ is not ramified in $L/K$. Passing to the completions at Weierstrass-$\wp$ put
@@
Because Weierstrass-$\wp$ does not divide $|G|$ the order $\Gamma_{\wp}$ is a maximal order in ${\mathcal{A}}$. As Weierstrass-$\wp$ does not divide $2\operatorname{disc}(L/K)$ we have ${\mathbb{Z}}_{L}\Delta_{\wp}=\Gamma_{\wp}$ and hence $\Delta_{\wp}$ is a maximal order in $A$. Both orders are invariant under the adjoint involution $\iota_{H}$. In particular for any $\Delta_{\wp}$-lattice $\Lambda\leq V_{K,\wp}$ also its $H$-dual lattice $\Lambda^{*}$ is $\Delta_{\wp}$-invariant. As $\Delta_{\wp}$ is a maximal order, there is $n\in{\mathbb{Z}}$ such that $\Lambda^{*}=\wp^{n}\Lambda$ and hence Weierstrass-$\wp$ is not ramified in the discriminant algebra of $H$. ∎
@@ 8.1 Orthogonal condensation
Condensation methods play an important role in the construction of (modular) character tables and decomposition matrices for large finite groups (see for instance [16] for a brief description of the general idea and one of the many papers citing this article for further applications).
Fixed point condensation is an important tool to compute orthogonal discriminants [3, Section 3.3.2] for large degree characters due to the availability of very sophisticated programs to compute in large permutation representations (in particular in [21]). Let $W\cong{\mathbb{Z}}^{[G:U]}$ and $\rho_{W}:G\to\operatorname{GL}(W)$ be such a permutation representation of the finite group $G$ on the cosets of the subgroup $U$. For $S\leq G$ put
@@
to denote the projection on the $S$-fixed points and
$N_{S}:=N_{G}(S)$
to be the normaliser of $S$ in $G$. Then $e_{S}$ is an idempotent in ${\mathbb{Z}}[\frac{1}{|S|}]G$ that is invariant under the natural involution $\iota$ of the group algebra with $\iota(g)=g^{-1}$ for all $g\in G$.
@@
Moreover $V_{2}$ is a $KN_{S}$-module and $Q_{2}$ is an $N_{S}$-invariant quadratic form on $V_{2}$. If this module $V_{2}$ is orthogonally stable as $N_{S}$-module, then we can predict the orthogonal discriminant of $(V_{2},Q_{2})$ and obtain the discriminant of $(V,Q)$ by computing the one of the condensed module $(V_{1},Q_{1})$. If the restriction of $V$ to $N_{S}$ is not orthogonally stable then also $(V_{1},Q_{1})$ is not an orthogonally stable $N_{S}$-module and there is no group that we can use to construct the invariant quadratic form $Q_{1}$. Thanks to [12, Proposition 2.2] (see Section 7) we can use the involution $\iota$ on $A^{\prime}$ to compute the discriminant of the polarisation of $Q_{1}$:
@@
If $K$ is a number field, then the computation of the composition factors of the condensed module $Ve_{S}$ is in general unfeasible as meat-axe methods only work well over finite fields. However, given an a-priori list of possible determinants of the composition factor $(V_{1},Q_{1})$ (e.g. all square classes $d(K^{\times})^{2}$ for which $K[\sqrt{d}]/K$ is only ramified at primes dividing the group order), then it is possible to deduce the correct square class by computing the determinant of $(V_{1},Q_{1})$ modulo enough well chosen prime ideals (that usually do not divide the group order).
@@ 8.2 Unitary condensation
The sporadic simple Harada-Norton group $HN$ has order $2^{14}\cdot 3^{6}\cdot 5^{6}\cdot 7\cdot 11\cdot 13$ and two pairs of complex conjugate indicator $o$ absolutely irreducible characters of even degree: $\chi_{25}=\overline{\chi_{26}}$ of degree $656250$ and character field ${\mathbb{Q}}(\chi_{25})={\mathbb{Q}}[\sqrt{-19}]$ and $\chi_{35}=\overline{\chi_{36}}$ of degree $1361920$ and character field ${\mathbb{Q}}(\chi_{35})={\mathbb{Q}}[\sqrt{-10}]$. All Schur indices of $HN:2$ are 1.
$\chi_{25}$ is irreducible modulo $5,7,19$, and has two odd degree constituents modulo $11$. As 19 is ramified in the character field, its 19-modular reduction is orthogonal, and from [1] we obtain $\operatorname{disc}(\chi_{25}\pmod{19})=O+$. As $5,7,$ and $11$ are norms the possible unitary discriminants of $\chi_{25}$ are $-2$ and $-3$.
Similarly $\chi_{35}$ is irreducible modulo $2,7,19$ and has two odd degree constituents modulo $11$. As $11$ is a norm in ${\mathbb{Q}}[\sqrt{-10}]/{\mathbb{Q}}$ the possible discriminants are 1,3,$5,1$ 3 , 5 , 15.
[…]
@@
8.4 An example: The group SU$(3,7)$.
For a smaller example consider the group
@@
$\begin{array}[]{|c|c|c|c|}\hline\cr\operatorname{{\mathfrak{X}}}&\operatorname{{\mathfrak{X}}}(1)&{\mathbb{Q}}(\operatorname{{\mathfrak{X}}})&\operatorname{disc}(\operatorname{{\mathfrak{X}}})\\ \hline\cr 13,14&258&{\mathbb{Q}}[\sqrt{-1}]&-7\\ \hline\cr 15,16&258&{\mathbb{Q}}[\sqrt{-2}]&-7\\ \hline\cr 27,28&344&{\mathbb{Q}}[\sqrt{-1}]&21\\ \hline\cr 29-32&344&{\mathbb{Q}}[\zeta_{8}]&7\\ \hline\cr 33-36&344&{\mathbb{Q}}[\zeta_{8}]&1\\ \hline\cr 37-44&344&{\mathbb{Q}}[y^{\prime\prime}_{48}]&1\\ \hline\cr 45-58&384&{\mathbb{Q}}[u_{43},\sqrt{-43}]&1\\ \hline\cr\end{array}$
A Sylow 7-subgroup $S$ of $G$ consists of the upper triangular matrices with 1 on the diagonal; the normaliser $B$ of $S$ is the group of all upper triangular matrices $B\cong 7^{1+2}:C_{48}$. The group $B$ has 48 linear characters, the ones that restrict trivially to $S$, one character, $\psi$, of degree $48$, restricting to the sum of the other 48 degree one characters of $S$, and $8$ characters $\chi_{1},\ldots,\chi_{8}$ of degree 42. $\chi_{1}$ has indicator --, $\chi_{2}$ indicator + and rational character field, $\chi_{3}=\overline{\chi_{4}}$ have character field ${\mathbb{Q}}(\sqrt{-1})$ and $\chi_{5},\ldots,\chi_{8}$ have character field ${\mathbb{Q}}(\zeta_{8})$. By Section 4.3 the unitary discriminant of $\chi_{1}$ over the relevant character fields is $-7$. The characters $\psi$ and $\chi_{2}$ is orthogonal, so [11, Corollary 4.4] implies that $\operatorname{disc}(\chi_{2})=-7$ and $\operatorname{disc}(\psi)=1$. We compute the remaining unitary discriminants by explicitly constructing the representation and an invariant Hermitian form over the character field. We obtain $\operatorname{disc}(\chi_{3})=\operatorname{disc}(\chi_{4})=-7$ and $\operatorname{disc}(\chi_{5})=\ldots=\operatorname{disc}(\chi_{8})=-1$.
The characters $\operatorname{{\mathfrak{X}}}_{13},\ldots,\operatorname{{\mathfrak{X}}}_{16}$ have a unitary stable restriction to $B$ of discriminant $-7$. The restriction to $B$ of $\operatorname{{\mathfrak{X}}}_{45},\ldots,\operatorname{{\mathfrak{X}}}_{58}$ is the sum $\psi+\chi_{1}+\ldots+\chi_{8}$ and hence unitary stable of discriminant $1$.
@@
$\begin{array}[]{|c|c|c|c|c|}\hline\cr\operatorname{{\mathfrak{X}}}_{V}&\operatorname{{\mathfrak{X}}}_{V(1-e_{S})}&\operatorname{disc}(V(1-e_{S}))&\operatorname{{\mathfrak{R}}}_{Ve_{S}}(t)&\operatorname{disc}(Ve_{S})\\ \hline\cr 27&\chi_{1}+\chi_{3,4}+\chi_{5-8}+\psi&-7&(\zeta_{12}^{-1},\zeta_{12}^{7})&-3\\ \hline\cr 28&\chi_{1}+\chi_{3,4}+\chi_{5-8}+\psi&-7&(-\zeta_{12}^{-1},-\zeta_{12}^{7})&-3\\ \hline\cr 29&\chi_{1}+\chi_{2}+\chi_{4}+\chi_{5-8}+\psi&-7&(-\zeta_{24}^{-5},-\zeta_{24}^{11})&-3\\ \hline\cr 30&\chi_{1}+\chi_{2}+\chi_{3}+\chi_{5-8}+\psi&-7&(\zeta_{24},\zeta_{24}^{-7})&-3\\ \hline\cr 31&\chi_{1}+\chi_{2}+\chi_{4}+\chi_{5-8}+\psi&-7&(\zeta_{24}^{-5},\zeta_{24}^{11})&-3\\ \hline\cr 32&\chi_{1}+\chi_{2}+\chi_{3}+\chi_{5-8}+\psi&-7&(-\zeta_{24},-\zeta_{24}^{-7})&-3\\ \hline\cr 33&\chi_{1-4}+\chi_{6}+\chi_{7}+\chi_{8}+\psi&-1&(-\zeta_{16},\zeta_{16})&-2-\sqrt{2}\\ \hline\cr 34&\chi_{1-4}+\chi_{5}+\chi_{6}+\chi_{7}+\psi&-1&(-\zeta_{16}^{7},\zeta_{16}^{7})&-2-\sqrt{2}\\ \hline\cr 35&\chi_{1-4}+\chi_{5}+\chi_{7}+\chi_{8}+\psi&-1&(-\zeta_{16}^{5},\zeta_{16}^{5})&-2+\sqrt{2}\\ \hline\cr 36&\chi_{1-4}+\chi_{5}+\chi_{6}+\chi_{8}+\psi&-1&(-\zeta_{16}^{3},\zeta_{16}^{3})&-2+\sqrt{2}\\ \hline\cr 37&\chi_{1-4}+\chi_{5}+\chi_{6}+\chi_{8}+\psi&-1&(\zeta_{48}^{-7},\zeta_{48})&-3\\ \hline\cr\end{array}$
All the characters $\operatorname{{\mathfrak{X}}}_{37-44}$ restrict to $B$ as the sum $\chi_{1}+\chi_{2}+\chi_{3}+\chi_{4}+\psi+$ three characters of $\{\chi_{5},\ldots,\chi_{8}\}$ and the sum of two faithful characters of $B/S$. This yields that the discriminant of $V(1-e_{S})$ is $-1$ for $\operatorname{{\mathfrak{X}}}_{37-44}$.

### v2

@@ 9.1 The fixed algebra under an automorphism
Let $G$ be a finite group and $\chi\in\operatorname{Irr}^{o}(G)$. Put $2m:=\chi(1)$, $L={\mathbb{Q}}(\chi)$, and assume that there is an $L$-representation $\rho:G\to\operatorname{GL}(V)$ affording the character $\chi$. Put $K:=L^{+}$ to denote the real subfield of $L$ and let $H:V\times V\to L$ be a $\rho(G)$-invariant $L/K$-Hermitian form on $V$. Assume that there is $\alpha\in\operatorname{Aut}(G)$ with $\alpha^{2}=1$ such that $\chi\circ\alpha=\overline{\chi}$. Then $\chi+\overline{\chi}$ extends to an irreducible character
$\operatorname{{\mathfrak{X}}}={\textrm{Ind}}_{G}^{\mathcal{G}}(\chi)\mbox{ of the semidirect product }{\mathcal{G}}:=G:\langle\alpha\rangle.$
@@
    Then $A]=[{\mathcal{Q}}]\in\operatorname{Br}_{2}(L,K)$.
@@
    Assume that the Frobenius-Schur indicator of $\operatorname{{\mathfrak{X}}}$ is ++.
    Then $A$ is an orthogonal subalgebra of $(\operatorname{End}_{L}(V),\iota_{H})$ and $\operatorname{disc}(H)=\operatorname{disc}_{L}([{\mathcal{Q}}])^{m}\operatorname{disc}(\iota_{|A})$.
@@
    Assume that the Frobenius-Schur indicator of $\operatorname{{\mathfrak{X}}}$ is -.
    Then $\operatorname{disc}(H)=\operatorname{disc}_{L}({\mathcal{Q}})^{m}$.
@@
To identify the class of the central simple $K$-algebra AA in the Brauer group, we relate the centraliser of $A$ in $\operatorname{End}_{K}(V)$ to the Schur indices of $\operatorname{{\mathfrak{X}}}$, i.e. to the class ${\mathcal{Q}}]\in\operatorname{Br}_{2}(L,K)$. We closely follow [2, Section 2]:
@@
Then $\overline{X}X$ commutes with all $\rho(g)$, so $\overline{X}X=\lambda I_{2m}$ for some $\lambda\in L^{\times}$. It is easy to see that $\lambda=\overline{\lambda}\in K^{\times}$ and is well defined up to norms. Now the proof of [2, Proposition 4.2] shows that ${\mathcal{Q}}]=[(L,\lambda)_{K}]$.
@@
(c) To apply Proposition 5.12 (a) it remains to show that the restriction of $\iota_{H}$ to $A$ is an orthogonal involution on $A$, i.e. that the dimension of skew-adjoint elements in $A$ is m(2m1)m(2m-1). This follows from the assumption on the indicator of $\operatorname{{\mathfrak{X}}}$. As dimensions remain unchanged under field extensions we may replace $K$ by the field $\mathbb{R}$} of real numbers and $L$ by $\mathbb{C}$}. The assumption on the Frobenius-Schur indicator of $\operatorname{{\mathfrak{X}}}$ shows that the Schur index of $\operatorname{{\mathfrak{X}}}$ over the real numbers $\mathbb{R}$} is 1. So the main result of [2] shows that there is a basis of $V\otimes_{L}{\mathbb{C}}$ such that the matrices in $\rho(G)$ with respect to this basis satisfy
@@
(d) Now assume that the indicator of $\operatorname{{\mathfrak{X}}}$ is -. Then by (a) the algebra $A\cong{\mathcal{Q}}^{m\times m}$ and the restriction of $\iota$ to $A$ is a symplectic involution. From Proposition 5.12 (b) one gets that $\Delta(H)=[{\mathcal{Q}}]^{m}$. ∎
@@
Let $\wp$ be a prime ideal of $K$ that does not divide $2|G|$. Then $\wp$ is not ramified in $L/K$. Passing to the completions at $\wp$ put
@@
Because $\wp$ does not divide $|G|$ the order $\Gamma_{\wp}$ is a maximal order in $\mathcal{A}$}. As $\wp$ does not divide $2\operatorname{disc}(L/K)$ we have ${\mathbb{Z}}_{L}\Delta_{\wp}=\Gamma_{\wp}$ and hence $\Delta_{\wp}$ is a maximal order in $A$. Both orders are invariant under the adjoint involution $\iota_{H}$. In particular for any $\Delta_{\wp}$-lattice $\Lambda\leq V_{K,\wp}$ also its $H$-dual lattice $\Lambda^{*}$ is $\Delta_{\wp}$-invariant. As $\Delta_{\wp}$ is a maximal order, there is $n\in{\mathbb{Z}}$ such that $\Lambda^{*}=\wp^{n}\Lambda$ and hence $\wp$ is not ramified in the discriminant algebra of $H$. ∎
@@ 10.1 Orthogonal condensation
Condensation methods play an important role in the construction of (modular) character tables and decomposition matrices for large finite groups (see for instance [19] for a brief description of the general idea and one of the many papers citing this article for further applications). Fixed point condensation is an important tool to compute orthogonal discriminants [3, Section 3.3.2] for large degree characters due to the availability of very sophisticated programs to compute in large permutation representations (in particular in [25]). Let $U\leq G$ and $W\cong{\mathbb{Z}}^{[G:U]}$ and $\rho_{W}:G\to\operatorname{GL}(W)$ be such a permutation representation of the finite group $G$ on the cosets of the subgroup $U$. For $S\leq G$ put
@@
to denote the projection on the $S$-fixed points. Then $e_{S}$ is an idempotent in ${\mathbb{Z}}[\frac{1}{|S|}]G$ that is invariant under the natural involution $\iota$ of the group algebra with $\iota(g)=g^{-1}$ for all $g\in G$.
@@
Let
$N_{S}:=N_{G}(S)$
be the normaliser of $S$ in $G$. Then $V_{2}$ is a $KN_{S}$-module and $Q_{2}$ is an $N_{S}$-invariant quadratic form on $V_{2}$. If this module $V_{2}$ is orthogonally stable as $N_{S}$-module, then we can predict the orthogonal discriminant of $(V_{2},Q_{2})$ and obtain the discriminant of $(V,Q)$ by computing the one of the condensed module $(V_{1},Q_{1})$. If the restriction of $V$ to $N_{S}$ is not orthogonally stable then also $(V_{1},Q_{1})$ is not an orthogonally stable $N_{S}$-module and there is no group that we can use to construct the invariant quadratic form $Q_{1}$. Thanks to [14, Proposition 2.2] (see Section 9) we can use the involution $\iota$ on $A^{\prime}$ to compute the discriminant of the polarisation of $Q_{1}$:
@@
If $K$ is a number field, then the computation of the composition factors of the condensed module $Ve_{S}$ is in general unfeasible as meat-axe methods only work well over finite fields. However, given an a priori list of possible determinants of the composition factor $(V_{1},Q_{1})$ (e.g. all square classes $d(K^{\times})^{2}$ for which $K[\sqrt{d}]/K$ is only ramified at primes dividing the group order), then it is possible to deduce the correct square class by computing the determinant of $(V_{1},Q_{1})$ modulo enough well chosen prime ideals (that usually do not divide the group order).
@@ 10.2 Unitary condensation
This example is intended to illustrate the power of condensation methods to handle large degree representations. We give some details that should allow the interesting reader to develop own condensation programs.
The sporadic simple Harada-Norton group $HN$ has order $2^{14}\cdot 3^{6}\cdot 5^{6}\cdot 7\cdot 11\cdot 13$ and two pairs of complex conjugate indicator $o$ irreducible characters of even degree: $\chi_{25}=\overline{\chi_{26}}$ of degree $656250$ and character field ${\mathbb{Q}}(\chi_{25})={\mathbb{Q}}[\sqrt{-19}]$ and $\chi_{35}=\overline{\chi_{36}}$ of degree $1361920$ and character field ${\mathbb{Q}}(\chi_{35})={\mathbb{Q}}[\sqrt{-10}]$. All Schur indices of $HN:2$ are 1.
$\chi_{25}$ is irreducible modulo $5,7,19$, and has two odd degree constituents modulo $11$. As 19 is ramified in the character field, its 19-modular reduction is orthogonal, and from [1] we obtain $\operatorname{disc}(\chi_{25}\pmod{19})=O+$. As $5,7,$ and $11$ are norms the possible unitary discriminants of $\chi_{25}$ are 2-2 and 3-3.
Similarly $\chi_{35}$ is irreducible modulo $2,7,19$ and has two odd degree constituents modulo $11$. As $11$ is a norm in ${\mathbb{Q}}[\sqrt{-10}]/{\mathbb{Q}}$ the possible discriminants are $1,3,5,15$.
The missing information is obtained using unitary condensation. The outer elements in $HN.2$ interchange $\chi_{25}$ and $\chi_{26}$ and also $\chi_{35}$ with $\chi_{36}$. The normaliser in $HN.2$ of the Sylow 5-subgroup does not contain an outer element of order $2$. So we take a subgroup $S$ of order $5^{5}$ whose normaliser $N_{S}:=N_{HN}(S)$ is the maximal subgroup of order $2000000$ of $HN$. Here we find an element $\alpha\in HN.2\setminus HN$ of order 2 that normalises $S$. The group $S$ has a $210$-dimensional fixed space on the module with character $\chi_{25}$ and a $416$-dimensional fixed space for $\chi_{35}$. We start with a permutation representation on the $16500000$ cosets of the maximal subgroup $U_{3}(8).3$ of $HN$. All four characters, $\chi_{25},\chi_{26},\chi_{35},$ and $\chi_{36}$ occur with multiplicity one in the corresponding permutation character. The group $S$ has $5280$ orbits, so we need to compute the action of a skew adjoint $\alpha$-fixed element on the composition factors of dimension $210$ respectively $416$ of a $5280$-dimensional module. We reduce modulo primes where 19-19 (resp. 10-10) are squares. It turns out that the $S$-fixed part of the $\alpha$-determinant of both, $\chi_{25}$ and $\chi_{35}$, is $33$. On the orthogonal complement, the normaliser $N_{S}$ acts unitary stably and we compute the discriminant here as $1$. As $11$ is a norm in both imaginary quadratic extensions we get
@@
10.4 An example: The group SU(3,7)(3,7).
[…]
@@
[…]
A Sylow 7-subgroup $S$ of $G$ consists of the upper triangular matrices with 1 on the diagonal; the normaliser $B$ of $S$ is the group of all upper triangular matrices $B\cong 7^{1+2}:C_{48}$. The group $B$ has 48 linear characters, the ones that restrict trivially to $S$, one character, $\psi$, of degree $48$, restricting to the sum of the other 48 degree one characters of $S$, and $8$ characters $\chi_{1},\ldots,\chi_{8}$ of degree 42. $\chi_{1}$ has indicator -, $\chi_{2}$ indicator + and rational character field, $\chi_{3}=\overline{\chi_{4}}$ have character field ${\mathbb{Q}}(\sqrt{-1})$ and $\chi_{5},\ldots,\chi_{8}$ have character field ${\mathbb{Q}}(\zeta_{8})$. By Proposition 5.12 (b) the unitary discriminant of $\chi_{1}$ over the relevant character fields is 7-7. The characters $\psi$ and $\chi_{2}$ are orthogonal, so [13, Corollary 4.4] implies that $\operatorname{disc}(\chi_{2})=-7$ and $\operatorname{disc}(\psi)=1$. We compute the remaining unitary discriminants by explicitly constructing the representation and an invariant Hermitian form over the character field. We obtain $\operatorname{disc}(\chi_{3})=\operatorname{disc}(\chi_{4})=-7$ and $\operatorname{disc}(\chi_{5})=\ldots=\operatorname{disc}(\chi_{8})=-1$.
The characters $\operatorname{{\mathfrak{X}}}_{13},\ldots,\operatorname{{\mathfrak{X}}}_{16}$ have a unitary stable restriction to $B$ of discriminant 7-7. The restriction to $B$ of $\operatorname{{\mathfrak{X}}}_{45},\ldots,\operatorname{{\mathfrak{X}}}_{58}$ is the sum $\psi+\chi_{1}+\ldots+\chi_{8}$ and hence unitary stable of discriminant $1$.
@@
[…]
All the characters $\operatorname{{\mathfrak{X}}}_{37-44}$ restrict to $B$ as the sum $\chi_{1}+\chi_{2}+\chi_{3}+\chi_{4}+\psi+$ three characters of $\{\chi_{5},\ldots,\chi_{8}\}$ and the sum of two faithful characters of $B/S$. This yields that the discriminant of $V(1-e_{S})$ is 1-1 for $\operatorname{{\mathfrak{X}}}_{37-44}$.